No. of solution of x (real) if [x] represents greatest integer function and {x} represents x – [x] :
7[x] + 23{x} = 191

Grade:12

2 Answers

Sher Mohammad IIT Delhi
9 years ago
7x-7{x}+23{x}=191
7x+16{x}=191
16{x}=191-7x
0<={x}<1
so 0<=191-7x<16

-16<7x-191<=0
175<7x<=191

175/7<x<=191/7

Sher Mohammad
IIT DELHI
faculty askiitians
Akshi
9 years ago
Sir, if I put x=26, it isn’t satisfying the eq. Moreover putting any x>175/7 and <191/7 isn’t satisfying the eq! Plz help.